OpenStudy (anonymous):

4z + b = 2z + c, solve for z

jhonyy9 (jhonyy9):

like a first step subtract from both sides c see what will get

OpenStudy (anonymous):

What?

jhonyy9 (jhonyy9):

the same like here 4z + b = 2z + c / - c 4z +b -c = 2z +c -c continue and write it here what will get

OpenStudy (anonymous):

4z + b = 2z + c 2z + b = c 2z = c - b z = c-b/2

OpenStudy (anonymous):

Dude your pic is hilarious.

OpenStudy (anonymous):

Anyway, how is the answer a fraction?

OpenStudy (anonymous):

haha thanks! What do you mean? We don't know what c or b is and we can't find out from what we're given so it has to be

OpenStudy (anonymous):

Oooooh. Okay thanks!

OpenStudy (anonymous):

No problem!